OpenStudy (anonymous):

write the standard form or the equation of the line parallel to x-2y-6=0 and passing through (0,1)

OpenStudy (amistre64):

dont delete your posts please ....

OpenStudy (amistre64):

parallel lines have the same equations; except for the constant is changed

OpenStudy (amistre64):

x-2y-c = 0 ; just plug in the values for x and y from the point they give youto calibrate it for the new "c"

OpenStudy (anonymous):

line parallel to the given line will have the same slope.. therefore m = 1/2 y=mx+b 1 = 1/2(0) + b => b= 1 Line is y = 1/2 x + 1